OpenStudy (anonymous):

Jambalaya is a Cajun dish made from chicken, sausage, and rice. Simone is making a large pot of jambalaya for a party. Chicken costs $6 per pound, sausage costs $3 per pound, and rice costs $1 per pound. She spends $42 on 13 and a half pounds of food. She buys twice as much rice as sausage. Using a system of equations, how much sausage will Simone use in her dish? a.4.5 lb of sausage c.6 lb of sausage b.3 lb of sausage d.12 lb of sausage

OpenStudy (anonymous):

How many pounds of food is it?

OpenStudy (anonymous):

i fixed it lol but its 13 and a half

OpenStudy (anonymous):

@AnonymousBeast

OpenStudy (anonymous):

@mathmate can you help with this one? i have no idea how to start it.

OpenStudy (mathmate):

c=chicken,s=saussage,r=rice ok?

OpenStudy (mathmate):

what is c+s+r?

OpenStudy (anonymous):

i think i figured it out thats why i closed it. but thank you very much again. (:

OpenStudy (anonymous):

I can set up the system, I just cannot remember how to solve it. Sorry :/

OpenStudy (mathmate):

So what did you get for your solution?

OpenStudy (mathmate):

@alicatforever So have you figured it out, and what is the problem, if there is any?

OpenStudy (anonymous):

i cant remember if i put a or b but it was right

OpenStudy (mathmate):

In that case, I will try to see which one you got: c+s+r=13.5 since r=2s so c+3s=13.5 .....(1) (reduces 1 unknown) For the cost 6c+3s+1(2s)=$42 6c+5s=42 .....(2) Multiply (1) by 6 6c+18s=81 Subtract (2) 6c+5s=42 We get 13s=39 s=3 Substitute in 1 c+3(3)=13.5 c=13.5-9 =4.5 As you said, it's (A)

OpenStudy (mathmate):

Sorry, it was B, 3 lbs of saussage. We didn't even need to calculate chicken. lol